In the period of a year, an inventor lost twice as much as he gained. If he originally had $10,000 and now has $8000, what amount did he gain and what did he lose?
Answer by ankor@dixie-net.com(22740)

In the period of a year, an inventor lost twice as much as he gained. If he originally had $10,000 and now has $8000, what amount did he gain and what did he lose?
:
Let x = amt gained
Then
2x = amt lost
:
10000 + x - 2x = 8000
:
x - 2x = 8000 - 1000
:
-x = - 2000
:
x = +2000 he gained and 4000 he lost